# Beamforming Design for Intelligent Reflecting Surface Aided Near-Field THz Communications

### v1

@@ Abstract
Intelligent reflecting surface (IRS) operating in the terahertz (THz) band has recently gained considerable interest due to its high spectrum bandwidth. Due to the exploitation of large scale of IRS, there is a high probability that the transceivers will be situated within the near-field region of the IRS. Thus, the near-field beam split effect poses a major challenge for the design of wideband IRS beamforming, which causes the radiation beam to deviate from its intended location, leading to significant gain losses and limiting the efficient use of available bandwidths. While delay-based IRS has emerged as a potential solution, current beamforming schemes generally assume unbounded range time delays (TDs). In this letter, we first investigate the near-filed beam split issue at the IRS. Then, we extend the piece-wise far-field model to the IRS, based on which, a double-layer delta-delay (DLDD) IRS beamforming scheme is proposed. Specifically, we employ an element-grouping strategy and the TD imposed on each sub-surface of IRS is achieved by a series of TD modules. This method significantly reduces the required range of TDs. Numerical results show that the proposed DLDD IRS beamforming scheme can effectively mitigate the near-field beam split and achieve near-optimal performance.
@@
IRS, THz, near-field, and beam split.
@@ I Introduction
Terahertz (THz) communication is a promising technique for future sixth-generation (6G) wireless communications due to its ultra-wide bandwidth, which is expected to achieve data rates of up to terabits per second (Tb/s) theoretically [1]. However, THz signals face significant challenges, including severe transmission attenuation and poor scattering, which limits the transmission range. Fortunately, intelligent reflecting surface (IRS) is a promising solution to these challenges. By using a large number of passive reflection elements, an IRS can steer the signals to any desired direction [2, 3]. This capability allows the IRS to enhance signal strength, making it a practical solution for future THz communications [4].
@@
Motivated by the delta-delay-phase precoding proposed in [10] that effectively reduces the maximum range of TD modules, we propose a delay-based IRS beamforming to overcome the aforementioned near-field beam split challenge in this letter. By extending the piece-wise far-field model in [11] to the uniform planar array (UPA) structured IRS, we divide the IRS into multiple sub-surfaces, each of which is imposed by a same TD value, and reasonably assume that the BS and the user are in the far-field range of each small sub-surface. This partition enables us to decompose the phase difference among different IRS elements that causes beam split into two separate components, i.e., the inter-surface phase difference and the intra-surface phase difference. To alleviate the inter-surface phase discrepancy, we propose a double-layer delta-delay (DLDD) architecture. In this design, the first layer TD network takes care of the phase discrepancy along one axis of IRS, while the second layer takes care of that along the other axis. Numerical results reveal the necessity of delay-enabled frequency-dependent IRS beamforming and demonstrate the effectiveness of the proposed scheme in mitigating the near-field beam split effect. Furthermore, it shows that the proposed DLDD IRS beamforming scheme can achieve near-optimal performance using short-range practical TD modules.
@@ II System Model
We consider an IRS-aided THz communication system, where an IRS is deployed to assist the downlink transmission from the BS to a user. In this letter, we focus on the beam split issue at the IRS, thus the BS and the user are assumed equipped with single antenna. Both of them are assumed to be within the Fraunhofer distance of the IRS, i.e., $R=\frac{2D^{2}}{\lambda_{c}}$, where $D$ is the maximum aperture of the IRS and $\lambda_{c}$ denotes the corresponding signal wavelength. As shown in Fig. 1, we assume that the IRS is located on the $y-z$ plane and centered at the origin in the Cartesian coordinates, which has $N_{y}$ elements along the $y$-axis and $N_{z}$ elements along the $z$-axis, satisfying $N=N_{y}\times N_{z}$. The IRS elements are assumed to be separated by $d$. Then, the coordinate of the $(n_{y},n_{z})$-th IRS element, with $n_{y}=\{1,\dots,N_{y}\}$ and $n_{z}=\{1\dots,N_{z}\}$, is given by $(0,\Delta_{n_{y}}^{N_{y}}d,\Delta_{n_{z}}^{N_{z}}d)$, where $\Delta_{a}^{b}=a-\frac{b-1}{2}$, with $a=\{0,\dots,b-1\}$ and $b=\{N_{y},N_{z}\}$. The coordinate of BS is $(x_{b},y_{b},z_{b})$ and that of user is $(x_{u},y_{u},z_{u})$. The center frequency is denoted by $f_{c}$ and the bandwidth is denoted by $B$, which is equally divided into $M$ subcarriers. Then, each subcarrier frequency can be denoted by $f_{m}=f_{c}+\frac{B}{M}\left(m-1-\frac{M-1}{2}\right)$, with $\{m=0,1,\dots,M-1\}$.
Assuming that the direct link between the BS and the user is blocked by obstructions, the user can only receive signals that are reflected by the IRS. In the THz band, the power gains associated with scattering paths are notably lower compared to the direct line-of-sight (LoS) path, thus we only focus on the LoS channel [7]. The LoS near-field channel between the BS and the IRS at the $m$-th subcarrier is denoted by $\mathbf{g}_{m}\in\mathbb{C}^{N\times 1}$, whose entries are given by
$\left[\mathbf{g}_{m}\right]_{n_{y},n_{z}}=\frac{\alpha_{m}}{r^{br}_{n_{y},n_{z}}}e^{-j2\pi\frac{f_{m}}{c}r^{br}_{n_{y},n_{z}}},\vspace{-0.2cm}$ (1)
where $\frac{\alpha_{m}}{r^{br}_{n_{y},n_{z}}}=\frac{c}{4\pi f_{m}r^{br}_{n_{y},n_{z}}}$ denotes the free-space path loss and $r^{br}_{n_{y},n_{z}}$ denotes the distance between the BS and the $(n_{y},n_{z})$-th IRS element, which can be written as $r^{br}_{n_{y},n_{z}}=\left(x_{b}^{2}+\left(y_{b}-\Delta_{N_{y}}^{n_{y}}d\right)^{2}+\left(z_{b}-\Delta_{N_{z}}^{n_{z}}d\right)^{2}\right)^{\frac{1}{2}}$.
@@
$\left[\mathbf{h}_{m}\right]_{n_{y},n_{z}}=\frac{\alpha_{m}}{r^{ru}_{n_{y},n_{z}}}e^{-j2\pi\frac{f_{m}}{c}r^{ru}_{n_{y},n_{z}}},\vspace{-0.2cm}$ (2)
where $r^{ru}_{n_{y},n_{z}}=\left(x_{u}^{2}+\left(y_{u}-\Delta_{N_{y}}^{n_{y}}d\right)^{2}+\left(z_{u}-\Delta_{N_{z}}^{n_{z}}d\right)^{2}\right)^{\frac{1}{2}}$ represents the distance between the $(n_{y},n_{z})$-th IRS element and the user. The IRS reflection coefficients matrix is expressed as
@@
where $\theta_{n}\in[0,2\pi),n=\{1,\dots,N\}$, is the phase shift of the $n$-th IRS reflecting element [2].
II-A Near-field Beam Split Effect at the IRS
We define the normalized array gain of the IRS at the $m$-th subcarrier as
$\eta\left(f_{m}\right)\!=\!\frac{1}{N}\!\left|\sum_{n_{y}=1}^{N_{y}}\sum_{n_{z}=1}^{N_{z}}\!e^{-j2\pi\frac{f_{m}}{c}\left(r^{br}_{n_{y},n_{z}}\!-r^{ru}_{n_{y},n_{z}}\right)}\!e^{j\theta_{n_{y},n_{z}}}\right|.$ (4)
@@ II-A Near-field Beam Split Effect at the IRS
$\theta_{n_{y},n_{z}}=\frac{2\pi}{\lambda_{c}}\left(r^{br}_{n_{y},n_{z}}-r^{ru}_{n_{y},n_{z}}\right),\forall n_{y},n_{z}.$ (5)
Owing to the frequency-independent characteristic of (5), the generated beams cannot perfectly point towards the target location in the whole bandwidth, which leads to severe gain loss. In Fig. 2LABEL:sub@beamsplit, we illustrate this by plotting the normalized array gain at $f_{c}$ and two edge subcarriers, $f_{1}$ and $f_{M}$, achieved by (5), under the same setup that will be specified in Section IV. It can be seen that the near-field beam split causes the beams at different subcarriers to split towards different locations. This restricts the user to receive signals that are only around the center frequency.
@@
II-B Piece-wise Far-Field Channel Modeling for UPA IRS
@@
We group the adjacent elements of the IRS into a sub-surface that will be imposed by one common TD value. It is assumed that each sub-surface contains $S\times S$ elements, satisfying $S=N_{y}/K_{y}=N_{z}/K_{z}$. Thus, the number of sub-surfaces is $K=K_{y}\times K_{z}$. The coordinate of the center of $(k_{y},k_{z})$-th sub-surface, with $k_{y}=\{1,\dots,K_{y}\}$ and $k_{z}=\{1,\dots,K_{z}\}$, is then given as $\left(0,\Delta_{k_{y}}^{K_{y}}Sd,\Delta_{k_{z}}^{K_{z}}Sd\right)$. Inspired by the piece-wise far-field channel model in [11], we next derive the channel model for the considered UPA-structured IRS. Specifically, by dividing the entire IRS into multiple sub-surfaces, each of which contains much fewer elements compared to the original surface. This partition significantly reduces the near-field range for each sub-surface. As a result, even if the BS/user is within the near-field region of the entire IRS, it can be reasonably assumed to be in the far-field region of each individual sub-surface. Let $r_{br}$ be the distance between the center of IRS and the BS, which can be written as $r_{br}=\sqrt{x_{b}^{2}+y_{b}^{2}+z_{b}^{2}}$. Then, the distance between the center of $(k_{y},k_{z})$-th sub-surface can be written as
$r^{br}_{k_{y},k_{z}}=\sqrt{x_{b}^{2}+\left(y_{b}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}+\left(z_{b}-\Delta_{k_{z}}^{K_{z}}Sd\right)^{2}}$
$=$ $\left(r_{br}^{2}+\left(\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}-2\Delta_{k_{y}}^{K_{y}}Sdr_{br}\sin\left(\psi^{ele}_{br}\right)\sin\left(\psi^{azi}_{br}\right)\right.$
$\left.+\left(\Delta_{k_{z}}^{K_{z}}Sd\right)^{2}-2\Delta_{k_{z}}^{K_{z}}Sdr_{br}\cos\left(\psi^{ele}_{br}\right)\right)^{\frac{1}{2}}.$ (6)
@@
To this end, the distance $r^{k_{y},k_{z}}_{s_{y},s_{z}}$ between the $(s_{y},s_{z})$-th element, with $s_{y},s_{z}=\{1,\dots,S\}$, of the $(k_{y},k_{z})$-th sub-surface and the BS can be expressed as
$r^{br,k_{y},k_{z}}_{s_{y},s_{z}}=$ $\left(r^{br\text{ }2}_{k_{y},k_{z}}+\left(\Delta^{S}_{s_{z}}d\right)^{2}-2\Delta^{S}_{s_{z}}dr^{br}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{br,k_{y},k_{z}}\right)\right.$
$\left.+\left(\Delta^{S}_{s_{y}}d\right)^{2}-2\Delta^{S}_{s_{y}}dr^{br}_{k_{y},k_{z}}\sin\left(\psi^{ele}_{br,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{br,k_{y},k_{z}}\right)\right)^{\frac{1}{2}}$
$\stackrel{{\scriptstyle(a)}}{{\approx}}$ $r^{br}_{k_{y},k_{z}}-\Delta^{S}_{s_{z}}dr^{br}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{br,k_{y},k_{z}}\right)$
$-\Delta_{s_{y}}^{S}d\sin\left(\psi^{ele}_{br,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{br,k_{y},k_{z}}\right),$ (7)
where
$\vspace{-0.2cm}\sin\left(\psi^{azi}_{br,k_{y},k_{z}}\right)=\frac{y_{b}-\Delta_{k_{y}}^{K_{y}}Sd}{\left(x^{2}+\left(y_{b}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\right)^{\frac{1}{2}}},$ (8)
$\vspace{-0.2cm}\sin\left(\psi^{ele}_{br,k_{y},k_{z}}\right)=\frac{\left(x_{b}^{2}+\left(y_{b}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\right)^{\frac{1}{2}}}{r^{br}_{k_{y},k_{z}}},$ (9)
@@
$\vspace{-0.2cm}\cos\left(\psi^{ele}_{br,k_{y},k_{z}}\right)=\frac{z_{b}}{r^{br}_{k_{y},k_{z}}}.\vspace{0.2cm}$ (10)
@@
$(a)$ in (II-B) holds due to the first-order Taylor expansion $(1+x)^{\frac{1}{2}}\approx 1+\frac{1}{2}x$ and the ignorance of $\left(\Delta^{S}_{s_{y}}d\right)^{2}$ and $\left(\Delta^{S}_{s_{z}}d\right)^{2}$. Accordingly, the normalized near-field channel model between the BS and the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface is approximated as $\tilde{\mathbf{g}}_{m}$, whose entries are given by
$\left[\tilde{\mathbf{g}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{-j2\pi\frac{f_{m}}{c}r^{br}_{k_{y},k_{z}}}e^{j2\pi\frac{f_{m}}{c}\Delta^{S}_{s_{z}}dr^{br}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{br,k_{y},k_{z}}\right)}$
×$\times e^{j2\pi\frac{f_{m}}{c}\Delta_{s_{y}}^{S}d\sin\left(\psi^{ele}_{br,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{br,k_{y},k_{z}}\right)}.\vspace{-0.2cm}$ (11)
@@
Similarly, the normalized near-field channel between the IRS and the user can be approximated as $\tilde{\mathbf{h}}_{m}$, whose entries are given as
@@
$\left[\tilde{\mathbf{h}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{-j2\pi\frac{f_{m}}{c}r^{ru}_{k_{y},k_{z}}}e^{j2\pi\frac{f_{m}}{c}\Delta^{S}_{s_{z}}dr^{ru}_{k_{y},k_{z}}\cos\left(\psi^{ele}_{ru,k_{y},k_{z}}\right)}$
@@
×$\times e^{j2\pi\frac{f_{m}}{c}\Delta_{s_{y}}^{S}d\sin\left(\psi^{ele}_{ru,k_{y},k_{z}}\right)\sin\left(\psi^{azi}_{ru,k_{y},k_{z}}\right)},$ (12)
where the angles are similarly defined as in (8), (9) and (10).
III Proposed IRS Beamforming
In this section, we propose the DLDD IRS beamforming scheme, which enables frequency-dependent IRS beamforming using practical short-range TD modules
Fig. 3: The proposed DDLD IRS beamforming architecture.

### v2

@@ Abstract
Intelligent reflecting surfaces (IRSs) operating in the terahertz (THz) band have gained considerable interest due to their high spectral bandwidth. With large-scale IRS deployment, transceivers are likely to be in the near-field region, where the near-field beam split effect becomes a major challenge for wideband IRS beamforming. This effect causes beam deviation from the intended location, leading to significant gain loss and waste in spectrum resource. While delay-based IRS has emerged as a potential solution, current beamforming schemes generally assume unbounded range time delays (TDs). In this paper, we first investigate the near-filed beam split issue at the IRS. Then, we extend the piecewise far-field model to the IRS, based on which, a double-layer delta-delay (DLDD) IRS beamforming scheme is proposed. Specifically, we employ an element-grouping strategy and the TD imposed on each sub-surface of IRS is achieved by a series of TD modules. This method significantly reduces the required range of TDs.In addition, a sub-optimal joint beamforming scheme is proposed to address the double-stage beam split at the base station and the IRS. Numerical results show that the proposed schemes can effectively mitigate the near-field beam split and achieve near-optimal performance.
@@
IRS, THz, near-field, joint beamforming, and beam split.
@@ I Introduction
Terahertz (THz) communication is a promising technique for future sixth-generation (6G) wireless communications due to its ultra-wide bandwidth, which is expected to achieve data rates of up to terabits per second (Tb/s) theoretically [1]. However, THz signals face significant challenges, including severe transmission attenuation and poor scattering, which limits the transmission range. Fortunately, intelligent reflecting surface (IRS) is a promising solution to these challenges. By using a large number of passive reflection elements, an IRS can steer the signals to any desired direction [2]. This capability allows the IRS to enhance signal strength, making it a practical solution for future THz communications [3].
@@
Considering the above, we aim to investigate the mitigation of near-field beam split effect with practical short-range TD modules in an IRS-aided THz communication system in this paper. The main contributions are summarized as follows. First, we extend the piecewise far-field model in [10] to the uniform planar array (UPA)-structured IRS. By dividing the IRS into multiple sub-surfaces, the base station (BS) and the user can be reasonably assumed to be in the far-field range of each small sub-surface. Such segmentation enables us to decompose the phase difference among different IRS elements that causes beam split into two separate components, i.e., the inter-surface counterpart and the intra-surface counterpart. Second, motivated by the delta-delay-phase precoding proposed in [11], we propose a novel double-layer delta-delay (DLDD) architecture to alleviate the inter-surface phase discrepancy. Specifically, the first layer TD network takes care of the phase discrepancy along one axis of IRS, while the second layer takes care of that along the other axis. Third, the discussions are extended from single-antenna to multi-antenna system, where the double-stage beam split at both the BS and the IRS occurs. To combat with such effect, we propose a sub-optimal joint beamforming scheme. Finally, numerical results reveal the necessity of delay-enabled frequency-dependent IRS beamforming and demonstrate the effectiveness of the proposed DLDD architecture in mitigating the near-field beam split using practical short-range TD modules. Furthermore, it is shown that the proposed joint beamforming scheme achieves enhanced achievable rate performance with strong robustness in a wide bandwidth range.
@@ II System Model
We consider an IRS-aided THz communication system, where an IRS is deployed to assist the downlink transmission from the BS to a user.111 The design can be straightforwardly extended to the scenarios with multi-antenna user. Also, it can be applied to the scenario with the direct path by modifying the equivalent channel model accordingly. As shown in Fig. 1, suppose the BS is equipped with an $N_{t}$-element uniform linear array along the $z$-axis and the $N$-element UPA-structured IRS is located on the $y-z$ plane, both with antenna/elements separation of $d$. The center of BS is located at $(x_{\text{b}},y_{\text{b}},z_{\text{b}})$, and the coordinate of the $n_{t}$-th element is then given by $(x_{\text{b}},y_{\text{b}},z_{\text{b}}-\Delta_{n_{t}}^{N_{t}}d)$, where $\Delta_{a}^{b}=a-1-\frac{b-1}{2}$. The IRS is centered at the origin, with $N_{y}$ elements along the $y$-axis and $N_{z}$ elements along the $z$-axis, satisfying $N=N_{y}\times N_{z}$. Then, the coordinate of the $(n_{y},n_{z})$-th IRS element, with $n_{y}=\mathcal{N}_{y}\triangleq\{1,\dots,N_{y}\}$ and $n_{z}=\mathcal{N}_{z}\triangleq\{1\dots,N_{z}\}$, is given by $(0,\Delta_{n_{y}}^{N_{y}}d,\Delta_{n_{z}}^{N_{z}}d)$. The coordinate of the single-antenna user is assumed to be $(x_{\text{u}},y_{\text{u}},z_{\text{u}})$. Both of the BS and the user are considered to be within the Fraunhofer distance of the IRS, i.e., $R=\frac{2D^{2}}{\lambda_{c}}$, where $D$ is the maximum aperture of the IRS and $\lambda_{c}$ denotes the corresponding signal wavelength. The center frequency is denoted by $f_{c}$ and the bandwidth is denoted by $B$, which is equally divided into $M$ subcarriers. Then, each subcarrier frequency can be denoted by $f_{m}=f_{c}+\frac{B}{M}\left(m-1-\frac{M-1}{2}\right)$, with $m=\mathcal{M}\triangleq\{1,\dots,M\}$.
Assuming that the direct link between the BS and the user is blocked by obstructions, the user can only receive signals that are reflected by the IRS. In the THz band, the power gains associated with scattering paths are notably lower compared to the direct line-of-sight (LoS) path, thus we only focus on the LoS channel [7].222 It is assumed the channels are perfectly estimated using the pilot-assisted methods [12]. The LoS near-field channel between the BS and the IRS at the $m$-th subcarrier is denoted by $\mathbf{G}_{m}=\left[\operatorname{vec}\left(\mathbf{g}_{m}^{1}\right),\cdots,\operatorname{vec}\left(\mathbf{g}_{m}^{N_{t}}\right)\right]\in\mathbb{C}^{N\times N_{t}}$, where the entries of $\mathbf{g}_{m}^{n_{t}}$ are given by
$\left[\mathbf{g}_{m}^{n_{t}}\right]_{n_{y},n_{z}}=\frac{\alpha_{m}}{r^{n_{t}}_{n_{y},n_{z}}}e^{-j2\pi\frac{f_{m}}{c}r^{n_{t}}_{n_{y},n_{z}}},$ (1)
where $\alpha_{m}\triangleq\frac{c}{4\pi f_{m}}$, $\frac{\alpha_{m}}{r^{n_{t}}_{n_{y},n_{z}}}$ denotes the free-space path loss, and $r^{n_{t}}_{n_{y},n_{z}}$ denotes the distance between the $n_{t}$-th BS antenna and the $(n_{y},n_{z})$-th IRS element, which can be calculated as $r^{n_{t}}_{n_{y},n_{z}}=\left(x_{\text{b}}^{2}+\left(y_{\text{b}}-\Delta_{n_{y}}^{N_{y}}d\right)^{2}+\left(z_{\text{b}}-\Delta_{n_{z}}^{N_{z}}d\right)^{2}\right)^{\frac{1}{2}}$.
@@
$\vspace{-0.2cm}\left[\mathbf{h}_{m}\right]_{n_{y},n_{z}}=\frac{\alpha_{m}}{l_{n_{y},n_{z}}}e^{-j2\pi\frac{f_{m}}{c}l_{n_{y},n_{z}}},$ (2)
where $l_{n_{y},n_{z}}=\left(x_{\text{u}}^{2}+\left(y_{\text{u}}-\Delta_{n_{y}}^{N_{y}}d\right)^{2}+\left(z_{\text{u}}-\Delta_{n_{z}}^{N_{z}}d\right)^{2}\right)^{\frac{1}{2}}$ represents the distance between the $(n_{y},n_{z})$-th IRS element and the user. The IRS reflection coefficients matrix is expressed as
@@
where $\theta_{n}\in[0,2\pi),n\in\{1,\dots,N\}$, is the phase shift of the $n$-th IRS reflecting element [4]. Then, the achievable rate per subcarrier can be calculated as
$R_{m}=\log_{2}\left(1+\frac{P_{m}\left|\mathbf{h}_{m}^{T}\boldsymbol{\Theta}\mathbf{G}_{m}\mathbf{w}_{m}\right|^{2}}{\sigma_{m}^{2}}\right),$ (4)
where $P_{m}$ is the BS transmit power, $\mathbf{w}_{m}$ is the BS beamforming vector, and $\sigma_{m}^{2}$ is the noise power at the $m$-th subcarrier.
III Wideband IRS Beamforming Scheme
In this section, the near-field beam split effect at the IRS is analyzed. Then, the piecewise far-field channel model for UPA-structured IRS is derived, based on which, a wideband IRS beamforming scheme based on DLDD architecture is proposed to address the beam split problem. The number of BS antenna is reduced to $N_{t}=1$ to simplify expression, while the scenario with multi-antenna BS will be discussed in Section IV. Thus, the BS-IRS channel is simplified to $\mathbf{g}_{m}$.
III-A Near-field Beam Split Effect at the IRS
We define the array gain of the IRS at the $m$-th subcarrier as
$\eta\left(f_{m}\right)$ $\!=\!\left|\mathbf{h}_{m}^{T}\boldsymbol{\Theta}\mathbf{g}_{m}\right|$ (5a)
$\!=\!\left|\sum_{n_{y}=1}^{N_{y}}\sum_{n_{z}=1}^{N_{z}}\!e^{-j2\pi\frac{f_{m}}{c}\left(r_{n_{y},n_{z}}\!+l_{n_{y},n_{z}}\right)}\!e^{j\theta_{n_{y},n_{z}}}\right|.$ (5b)
@@ III-A Near-field Beam Split Effect at the IRS
$\theta_{n_{y},n_{z}}=\frac{2\pi}{\lambda_{c}}\left(r_{n_{y},n_{z}}+l_{n_{y},n_{z}}\right),n_{y}\in\mathcal{N}_{y},n_{z}\in\mathcal{N}_{z}.$ (6)
Owing to the frequency-independent characteristic of (6), the generated beams cannot perfectly point towards the target location in the whole bandwidth, which leads to severe gain loss. In Fig. 2LABEL:sub@beamsplit, we illustrate this by plotting the normalized array gain at $f_{c}$ and two edge subcarriers, $f_{1}$ and $f_{M}$, achieved by (6), under the same setup that will be specified in Section IV. It can be observed that the near-field beam split causes the beams at different subcarriers to split towards different locations. This restricts the user to receive signals that are only around the center frequency.
@@
III-B Piecewise Far-Field Channel Modeling for UPA IRS
@@
To reduce hardware cost, we adopt the sub-connected structure. Specifically, we divide the IRS into $K$ sub-surfaces, with $K_{y}$ sub-surfaces along the $y$-axis and $K_{z}$ sub-surfaces along the $z$-axis, i.e., $K=K_{y}\times K_{z}$. It is assumed that each sub-surface contains the adjacent $S\times S$ elements that share a common delay, satisfying $S=N_{y}/K_{y}=N_{z}/K_{z}$. The coordinate of the center of $(k_{y},k_{z})$-th sub-surface, with $k_{y}\in\mathcal{K}_{y}\triangleq\{1,\dots,K_{y}\}$ and $k_{z}\in\mathcal{K}_{z}\triangleq\{1,\dots,K_{z}\}$, is then expressed as $\left(0,\Delta_{k_{y}}^{K_{y}}Sd,\Delta_{k_{z}}^{K_{z}}Sd\right)$. Inspired by the piecewise far-field channel model in [10], we next derive the channel model for the considered UPA-structured IRS. Specifically, by dividing the entire IRS into multiple sub-surfaces, each of which contains much fewer elements compared to the original surface. This partition significantly reduces the near-field range for each sub-surface. As a result, even if the BS/user is within the near-field region of the entire IRS, it can be reasonably assumed to be in the far-field region of each individual sub-surface. Let $r$ be the distance between the center of IRS and the BS, which can be written as $r=\sqrt{x_{\text{b}}^{2}+y_{\text{b}}^{2}+z_{\text{b}}^{2}}$. Then, the distance between the BS and the center of $(k_{y},k_{z})$-th sub-surface can be written as
$r_{k_{y},k_{z}}=$ $\sqrt{x_{\text{b}}^{2}+\left(y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}+\left(z_{\text{b}}-\Delta_{k_{z}}^{K_{z}}Sd\right)^{2}}$ (7a)
$=$ $\left(r^{2}+\left(\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}-2\Delta_{k_{y}}^{K_{y}}Sdr\sin\left(\psi\right)\sin\left(\phi\right)\right.$
$\left.+\left(\Delta_{k_{z}}^{K_{z}}Sd\right)^{2}-2\Delta_{k_{z}}^{K_{z}}Sdr\cos\left(\psi\right)\right)^{\frac{1}{2}},$ (7b)
@@
where $\phi$ and $\psi$ are the azimuth and elevation angles of the BS with respect to (w.r.t.) the center of the IRS, respectively, as illustrated in Fig. 1. To this end, the distance $r^{k_{y},k_{z}}_{s_{y},s_{z}}$ between the BS and the $(s_{y},s_{z})$-th element, with $s_{y},s_{z}\in\mathcal{S}\triangleq\{1,\dots,S\}$, of the $(k_{y},k_{z})$-th sub-surface can be expressed as
[…]
$\left.+\left(\Delta^{S}_{s_{y}}d\right)^{2}-2\Delta^{S}_{s_{y}}dr_{k_{y},k_{z}}\sin\left(\psi_{k_{y},k_{z}}\right)\sin\left(\phi_{k_{y},k_{z}}\right)\right)^{\frac{1}{2}}$ (8a)
$\stackrel{{\scriptstyle(a)}}{{\approx}}\!r_{k_{y},k_{z}}\!\!-\!\Delta^{S}_{s_{z}}d\cos\!\left(\psi_{k_{y},k_{z}}\!\right)\!-\!\Delta_{s_{y}}^{S}d\sin\!\left(\psi_{k_{y},k_{z}}\!\right)\!\sin\!\left(\phi_{k_{y},k_{z}}\!\right)\!,$ (8b)
where $\phi_{k_{y},k_{z}}$ and $\psi_{k_{y},k_{z}}$ are the azimuth and elevation angles of the BS w.r.t. the center of $(k_{y},k_{z})$-th IRS sub-surface, respectively, satisfying
$\sin\left(\phi_{k_{y},k_{z}}\right)=\frac{y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd}{\left(x_{\text{b}}^{2}+\left(y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\right)^{\frac{1}{2}}},$ (9)
$\vspace{-0.2cm}\sin\left(\psi_{k_{y},k_{z}}\right)=\frac{\left(x_{\text{b}}^{2}+\left(y_{\text{b}}-\Delta_{k_{y}}^{K_{y}}Sd\right)^{2}\right)^{\frac{1}{2}}}{r_{k_{y},k_{z}}},$ (10)
@@
$\vspace{-0.2cm}\cos\left(\psi_{k_{y},k_{z}}\right)=\frac{z_{\text{b}}-\Delta_{k_{z}}^{K_{z}}Sd}{r_{k_{y},k_{z}}}.\vspace{0.2cm}$ (11)
@@
$(a)$ in (8) holds due to the first-order Taylor expansion $(1+x)^{\frac{1}{2}}\approx 1+\frac{1}{2}x$ and the ignorance of $\left(\Delta^{S}_{s_{y}}d\right)^{2}$ and $\left(\Delta^{S}_{s_{z}}d\right)^{2}$. [^3] Such ignorance arises from the fact that the distance between the $r_{k_{y},k_{z}}$ is generally much larger than the element separations, $\Delta^{S}_{s_{y}}d$ and $\Delta^{S}_{s_{z}}d$, within a sub-surface. Accordingly, the normalized near-field channel model between the BS and the $(s_{y},s_{z})$-th IRS element in the $(k_{y},k_{z})$-th sub-surface is approximated as $\tilde{\mathbf{g}}_{m}$, whose entries are given by
$\vspace{-2cm}\left[\tilde{\mathbf{g}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{-j2\pi\frac{f_{m}}{c}r_{k_{y},k_{z}}}e^{j2\pi\frac{f_{m}}{c}\Delta^{S}_{s_{z}}d\cos\left(\psi_{k_{y},k_{z}}\right)}$
×$\times e^{j2\pi\frac{f_{m}}{c}\Delta_{s_{y}}^{S}d\sin\left(\psi_{k_{y},k_{z}}\right)\sin\left(\phi_{k_{y},k_{z}}\right)}.\vspace{-0.5cm}$ (12)
@@
Similarly, the normalized near-field channel between the IRS and the user can be approximated as $\tilde{\mathbf{h}}_{m}$, whose entries are given by
@@
$\vspace{-0.3cm}\left[\tilde{\mathbf{h}}_{m}\right]^{k_{y},k_{z}}_{s_{y},s_{z}}=$ $e^{-j2\pi\frac{f_{m}}{c}l_{k_{y},k_{z}}}e^{j2\pi\frac{f_{m}}{c}\Delta^{S}_{s_{z}}d\cos\left(\upsilon_{k_{y},k_{z}}\right)}$
@@
×$\times e^{j2\pi\frac{f_{m}}{c}\Delta_{s_{y}}^{S}d\sin\left(\upsilon_{k_{y},k_{z}}\right)\sin\left(\omega_{k_{y},k_{z}}\right)},\vspace{-0.2cm}$ (13)
where $l_{k_{y},k_{z}}$ is the distance between the user and the center of $(k_{y},k_{z})$-th sub-surface, while $\omega_{k_{y},k_{z}}$ and $\upsilon_{k_{y},k_{z}}$ are the corresponding azimuth and elevation angles, respectively, defined similarly as (9), (10) and (11).
III-C Proposed DLDD IRS Beamforming
Fig. 3: The proposed DLDD IRS beamforming architecture.